# FLARE v2: A Recursive Framework for Program Comprehension Across Languages and Levels of Abstraction

### v1

@@ 4.3 The recursive element-binding-element cycle
This cycle can be applied repeatedly. At one level, it explains how blocks bind into a segment; at the next, how segments bind into a system. Nothing in the cycle depends on visual form: it applies equally to block-based, text-based, and hybrid environments.
In classroom practice, the cycle is enacted through question stems rather than formal steps. A teacher might first stabilise the elements in view (”Which bits belong together?”) before moving to questions about what each receives or sends, then to when and under what conditions interactions occur, and finally to naming the new unit that emerges. Section 5 illustrates this on a concrete program.
It is important to distinguish FLARE v2’s formal grammar from its classroom realisation. The analytic labels - Causal-Temporal, Communicative, Sequential, Branch, Event - are not intended as classroom vocabulary. In practice, these appear as simple, scalable questions such as ”What happens next?”, ”What triggers this?”, ”Under what conditions?”, and ”What does this share?” The cognitive load reported in FLARE v1 arose partly from presenting a multi-tier structure all at once. FLARE v2 supports progressive reveal: the same family of questions can be rehearsed first on blocks, then on segments, and then on systems, without introducing new structural categories at each stage.
@@ 4.4 Relationship to existing program comprehension models
FLARE v2 can be situated relative to three established traditions in program comprehension research. It aligns with Pennington’s program-versus-situation model, Schulte’s Block Model, and Sorva’s account of notional machines, while extending each to event-driven environments typical of contemporary school curricula.
@@
Pennington showed that programmers construct two complementary mental representations: a program model that foregrounds control flow, and a situation model that foregrounds data flow and goals (pennington1987comprehension). FLARE v2 mirrors this duality in its two binding dimensions: Causal-Temporal binding captures how control moves, while Communicative binding captures how information and shared state move between elements. The main extension is to make event-driven control an explicit subtype, acknowledging hidden runtime mechanisms that Pennington’s procedural studies did not need to model.
@@
The Block Model conceptualises program comprehension as reasoning about multiple hierarchical levels (from atoms to macrostructure) along three dimensions: text surface, execution, and function (schulte2008block). FLARE v2 preserves this insight but refactors it: structure and flow are captured by the two binding dimensions, while function (purpose) is treated as an interpretive stance that can be applied at any scale. The key addition is the recursive element-binding-element cycle, which explains how meaning propagates between levels rather than treating levels as fixed.
@@
Sorva reviews evidence that learners form mental models of program execution whether or not explicit notional machines are taught, and that hidden runtime mechanisms are a major source of misconception in many environments (sorva2013notional). FLARE v2 responds by making Event binding explicit as part of the Causal-Temporal dimension and by separating visible and hidden control mechanisms. This gives teachers a way to surface the invisible event loops and schedulers that govern much of the block-based code children encounter, while remaining agnostic about the particular notional machine adopted.
Taken together, these correspondences position FLARE v2 as an integrative extension of classical procedural comprehension models to the event-driven, often concurrent artefacts used in school computing.
@@ 5 Illustrative Example: A Simple Control System
This section illustrates the element-binding-element cycle and the four properties using a minimal three-segment control system commonly encountered in introductory physical computing and robotics. The system, which we refer to as AutoDimmer, comprises a ReadSensor segment that samples the current light level, a DecideBrightness segment that determines an appropriate LED brightness based on a threshold, and a SetLED segment that applies the chosen brightness to the hardware. This sense-decide-act pattern is familiar in many classroom microcontroller projects. The example is intentionally simple and language-agnostic: its purpose is to make the element-binding-element cycle concrete rather than to present a complex or novel control design.
@@
AutoDimmer control system, illustrated here before the formal analysis in Table 4.
@@ 5.2 Step 2: Describe each segment with the four properties
Each segment can then be described using Receives, Sends, Effects, and Shares. Table 4 summarises this for the three segments.
At this stage, a teacher’s questions might be:
  • ”What does ReadSensor receive? What does it send? Does it change anything inside itself?”
  • ”What does DecideBrightness depend on? What shared values does it read?”
  • ”What does SetLED change in the world? Does anything else share that device?”
These questions apply the four properties to each segment without yet considering how the segments are connected.
@@ 5.3 Step 3: Analyse bindings between segments
The next step is to analyse how the segments bind together along the two binding dimensions. The same three segments illustrate both dimensions and multiple Causal-Temporal subtypes.
@@
The three segments exhibit different Causal-Temporal subtypes:
  • Sequential binding connects the segments within each cycle: ReadSensor completes, then DecideBrightness runs, then SetLED executes. No branch instruction governs these transitions; each segment simply follows the previous one.
  • Branch binding operates within DecideBrightness, where a conditional compares the sensor reading with the threshold and selects between alternative brightness outputs. The if/else construct creates a branch in the user’s visible code.
  • Event binding (or a timer-based equivalent) governs when the entire sense-decide-act cycle begins. If a forever loop or timer tick triggers each cycle, this trigger mechanism is often partially hidden - the learner writes ”forever” but may not see the underlying loop mechanics. In a purely event-driven version, a sensor-change event might trigger the cycle directly.
Typical questions for this dimension include:
  • ”What has to happen before DecideBrightness can do its job?” (Sequential)
  • ”Under what conditions does the LED get set to maximum?” (Branch)
  • ”What triggers this whole chain to start each time?” (Event or timer)
@@
The elements are linked by the values they exchange and the state they share:
  • ReadSensor sends the sampled light level; DecideBrightness receives it.
  • DecideBrightness sends a chosen brightness value; SetLED receives it.
  • The threshold value that DecideBrightness reads is a shared configuration that other elements in the wider program could also access or update.
  • The LED device that SetLED writes to is shared hardware - other subsystems might also control it.
Questions for this dimension include:
  • ”Which segments rely on, or pass along, the light level value?”
  • ”Where does the threshold come from, and who else can change it?”
  • ”What would happen if another part of the program also tried to control the LED?”
Note that the internal calculations within DecideBrightness - any temporary variables used to compute the brightness-are Effects and do not participate in bindings. They matter for understanding what happens inside the segment, but they create no relationships with other elements.
Taken together, these bindings show why the three segments form a coherent control subsystem rather than three unrelated actions. Both binding dimensions operate concurrently. The same conditional that creates a Branch binding (comparing light level to threshold) also depends on Communicative binding (reading the shared threshold and the received light level). In classroom talk, a teacher might foreground one dimension at a time, typically starting with Sequential flow, then layering in questions about conditions (Branch) and shared state (Communicative).
@@ 5.4 Step 4: Recognise the system as a new element
Once the bindings are understood, the three segments can be treated as a single element at the system layer: the AutoDimmer system. From this point on, the system can be described using the same four properties:
  • Receives: ambient light readings (from the sensor) and periodic triggers from the control loop or timer.
  • Sends: brightness commands addressed to the LED hardware.
  • Effects: optional internal bookkeeping such as smoothing counters or mode flags. In this minimal example, the system has no private internal state.
  • Shares: external resources it interacts with, including the LED device and the shared threshold used by the decision logic.
The physical change in LED brightness is not an internal Effect of the AutoDimmer system. It is a change to an external device and therefore a Share. The system sends a brightness command; the resulting change in device state is a write to shared external state.
Recognising the system as a new element completes one turn of the recursive cycle: blocks combine into segments; segments combine into a system; and the resulting system can now participate in further bindings at a higher scale.
5.5 Stepping up the ladder: system within a larger service
The same cycle applies at higher scales. In a typical microcontroller project, AutoDimmer might sit within a larger room-controller service alongside subsystems such as HeatControl and PresenceDetect. At this scale, elements are whole systems rather than individual segments.
A room-controller service might:
  • Receive sensor data from light, temperature, and motion detectors, plus external commands from a user interface or network.
  • Send commands to lighting, heating, and display devices, or outbound messages to a cloud service.
  • Effect internal scheduling or mode flags that are private to the service.
  • Share configuration, occupancy state, or hardware devices that other subsystems may also read or update.
The bindings between subsystems follow the same two dimensions:
  • Causal-Temporal: Which subsystem runs first? What events trigger each? Do any subsystems wait for others to complete (Sequential), respond to conditions set by others (Branch), or activate independently on external stimuli (Event)?
  • Communicative: What data flows between subsystems? What configuration or hardware state do they share?
Changes to heaters, LEDs, displays, or network endpoints are all writes to shared state, not internal Effects. The definitions of Receives, Sends, Effects, and Shares therefore remain stable as the recursion climbs the ladder.
@@ 6 Discussion: Pedagogical and Curricular Implications
The Compositional Ladder suggests a natural pattern of progression: learners first gain fluency with smaller elements and simple bindings, then encounter larger elements whose internal structure can be analysed using the same questions. This aligns with established principles of spiral curricula in which ideas are revisited with increasing sophistication. Rather than introducing a full hierarchy at once, teachers can dwell at a particular rung until learners are fluent with its elements and bindings, and then move upward when appropriate.
FLARE may be understood both as a descriptive grammar for analysing program structure and as a mental framework that may shape how teachers scaffold the construction of code. By providing a stable set of properties (Receives, Sends, Effects, Shares) and binding dimensions (Causal-Temporal with its three subtypes, and Communicative), FLARE offers novices a way to articulate what a segment is intended to do before or while constructing it. To be clear: FLARE v2 is not a psychological model of novice cognition. It is a semiotic-compositional account of program meaning intended to structure teacher questioning and learner reasoning.
The separation of Causal-Temporal subtypes may prove particularly valuable for event-driven environments. The Year 5 probe reported in Section 3 found that most children selected concurrent execution when a second event was triggered. Given that their only prior exposure had been a brief event-handling metaphor two years earlier, with no reinforcement since, this pattern suggests that early metaphors can support persistent concurrent execution models in event-driven environments.
Because the ladder parallels familiar literacy progressions, FLARE may reduce cognitive load for teachers whose professional expertise lies in literacy, though this remains to be tested empirically.
From a pedagogical perspective, PRIMM (sentance2019primm) offers a temporal sequence for classroom activity (Predict, Run, Investigate, Modify, Make). FLARE may complement this. PRIMM structures what learners do over time; FLARE provides a way to talk about what they see and think at each phase.
A simple illustration shows how the two can interact. During Predict, a teacher might ask pupils which elements they expect to be involved in a behaviour and what triggers them (”Which segments will run when the button is clicked?”-an Event binding question). In Run, these elements are observed and named. During Investigate, FLARE questions focus attention on Causal-Temporal flow (”What happens next? What conditions control this branch?”) and Communicative connections (”What does this segment share?”). In Modify and Make, pupils reuse the same vocabulary to assemble known elements into new arrangements (”What will this new segment receive and send?”). On this view, PRIMM provides the classroom rhythm, while FLARE supplies the underlying grammar that helps that rhythm culminate in genuinely compositional, creative work. These relationships remain speculative and require empirical evaluation.
@@
In curriculum design, the recursive model suggests a spiral structure in which new abstraction levels appear once earlier bindings are secure. Early years may emphasise Sequential flow within small elements; later years can introduce Branch bindings with their associated conditional reasoning, and then Event bindings with their hidden control mechanisms. This mirrors literacy development and may support teachers in extending familiar practices from reading and writing text to reading and ”writing” code structure. Because Macro is treated as an interpretive stance rather than a separate tier, questions of purpose and audience (”What is this element for?”, ”Who relies on it?”) can be introduced at any rung without adding another structural category.
@@ 7 Limitations and Future Work
This paper is conceptual and exploratory. It does not report new empirical data beyond the exploratory Year 5 probe reported in Section 3 and detailed in heath2025concurrency. That probe requires systematic replication with controlled conditions. The analysis is interpretive and practitioner-led, drawing on reflective re-reading of earlier classroom work rather than on a newly collected or systematically coded dataset.
The two-dimension binding scheme with three Causal-Temporal subtypes has been developed for imperative and event-driven contexts typical of school curricula. Its applicability to object-oriented, concurrent, or distributed paradigms remains untested. The claim that Event binding is ”mechanically equivalent to Branch but pedagogically distinct” requires further investigation: do learners actually benefit from this distinction, or does it introduce unnecessary complexity?
Future research should examine how the recursive binding cycle manifests in teacher and learner discourse and whether it transfers across languages, environments, and programming paradigms. Studies might analyse classroom dialogue using the element–binding–element cycle and the two binding dimensions, investigate how teachers use the Compositional Ladder to structure tasks and questions, or explore whether making Event binding explicit helps learners develop more accurate notional machines for event-driven code. Implementation studies and evaluations of classroom uptake lie beyond the present scope but will be essential for assessing the practical value of FLARE v2.
@@ 8 Conclusion
The shift from a four-tier descriptive hierarchy (Blocks, Segments, Relations, Macro) to a recursive element–binding cycle represents a move from classification toward generative explanation. FLARE v2 positions program comprehension as a semiotic process: bounded elements carry meaning, and new meaning emerges as elements bind through Causal-Temporal relations (Sequential, Branch, or Event) and Communicative relations (what flows between elements and what they share). This reframing provides a unified conceptual vocabulary for discussing structure, flow, and purpose across artefact scales.
The two-dimension scheme aligns with Pennington’s established distinction between control-flow and data-flow abstractions. Although Pennington’s original work concerned sequential procedural programs, the control-flow / data-flow distinction can be generalised to event-driven environments where control mechanisms are hidden from the learner. By treating Event as a distinct subtype of Causal-Temporal binding, FLARE v2 acknowledges both the mechanical reality (events are dispatched by branches in a runtime loop) and the pedagogical reality (learners cannot see that loop and may develop intuitions that differ from sequential-first assumptions).
FLARE v2 emerged from reflective classroom experience and is presented here as a theoretical proposal rather than an empirical claim. In computing education, practitioner-informed conceptual work serves an important role: it can articulate implicit classroom reasoning, connect practice with broader theoretical traditions, and motivate targeted empirical studies. The framework introduced here invites such investigation. Future work should examine how the element–binding cycle can support teacher questioning, pupil reasoning, and code creation; how it complements existing pedagogical approaches such as PRIMM; and how it may inform curriculum design across primary and secondary phases.
In this sense, the purpose of FLARE v2 is not to settle questions of program comprehension, but to open them. By offering a stable, scalable vocabulary grounded in both semiotic theory and classroom need, it aims to provide a basis for subsequent empirical evaluation, refinement, and dialogue within the computing education community.
@@ 11 Conflict of Interest
This preprint is made available under the Creative Commons Attribution–NonCommercial–ShareAlike 4.0 International Licence (CC BY-NC-SA 4.0). This permits copying, distribution, and adaptation for non-commercial purposes, provided that appropriate credit is given and that derivative works are released under the same licence.
This licence applies only to the text, figures, and explanations contained in this paper. Practical implementations of the framework—including training materials, classroom scaffolds, exemplar resources, and software tools—are licensed separately and are not covered by this CC BY-NC-SA 4.0 licence.

### v2

@@ 5.3 The recursive element-binding-element cycle
This cycle can be applied repeatedly. At one level, it explains how blocks bind into a segment; at the next, how segments bind into a system. In classroom practice, the cycle is enacted through question stems rather than formal steps.
@@ 5.4 Relationship to existing program comprehension models
FLARE v2 can be situated relative to three established traditions in program comprehension research. It aligns with Pennington’s program-versus-situation model, Schulte’s Block Model, and Sorva’s account of notional machines, while extending each to event-driven environments common in school curricula.
@@
Pennington showed that programmers construct two complementary mental representations: a program model that foregrounds control flow, and a situation model that foregrounds data flow and goals (pennington1987comprehension). FLARE v2 mirrors this duality in its two binding dimensions: Causal-Temporal binding captures how control moves, while Communicative binding captures how information and shared state move between elements. The extension here is to treat event-driven control as a distinct subtype because its dispatch mechanism is typically hidden from the learner.
@@
The Block Model conceptualises program comprehension as reasoning about multiple hierarchical levels (from atoms to macrostructure) along three dimensions: text surface, execution, and function (schulte2008block). FLARE v2 preserves the multi-level insight but makes the level-to-level transition explicit via the recursive element-binding-element cycle, rather than treating levels as a fixed descriptive ladder.
@@
Sorva reviews evidence that learners form mental models of program execution whether or not explicit notional machines are taught, and that hidden runtime mechanisms are a major source of misconception (sorva2013notional). FLARE v2 responds by separating visible and hidden control mechanisms within Causal-Temporal binding and by providing a descriptive vocabulary for surfacing hidden dispatch or scheduling in event-driven code.
@@ 6 Illustrative Example: A Simple Control System
This section illustrates the element-binding-element cycle and the four properties using a minimal three-segment control system commonly encountered in introductory physical computing and robotics. The system, which we refer to as AutoDimmer, comprises a ReadSensor segment that samples the current light level, a DecideBrightness segment that determines an appropriate LED brightness based on a threshold, and a SetLED segment that applies the chosen brightness to the hardware. The example is intentionally simple and language-agnostic: its purpose is to make the element-binding-element cycle concrete.
@@
AutoDimmer control system.
@@ 6.2 Step 2: Describe each segment with the four properties
Each segment can be described using Receives, Sends, Effects, and Shares (Table 4). At this stage, questions might include: ”What does ReadSensor receive and send?”, ”What does DecideBrightness read that is shared?”, and ”What does SetLED write to outside the code?”.
@@ 6.3 Step 3: Analyse bindings between segments
The next step is to analyse how the segments bind along the two dimensions.
@@
Sequential binding connects the segments within each cycle: ReadSensor runs, then DecideBrightness, then SetLED. Branch binding operates within DecideBrightness, where a conditional compares the sensor reading with the threshold and selects between alternative outputs. The system may also be governed by a timer tick or a loop trigger; depending on environment and representation, that trigger mechanism may be partially hidden from the learner, motivating Event-style questions (”What triggers this whole chain to start each time?”).
@@
ReadSensor sends the sampled light level and DecideBrightness receives it. DecideBrightness sends a brightness value and SetLED receives it. The threshold value is shared configuration, and the LED is shared hardware state. Any temporary calculations inside DecideBrightness are Effects and do not create inter-element bindings, consistent with the definition in Table 2.
@@ 6.4 Step 4: Recognise the system as a new element
Once the bindings are understood, the three segments can be treated as a single element at the system layer: the AutoDimmer system. From this point on, the system can be described using the same four properties (Receives, Sends, Effects, Shares), enabling the cycle to repeat at higher scale.
@@ 7 Discussion: Pedagogical and Curricular Implications
The Compositional Ladder suggests a progression in which learners first gain fluency with smaller elements and simple bindings, then encounter larger elements whose internal structure can be analysed using the same questions. This is compatible with spiral curriculum principles: concepts are revisited with increasing sophistication, but without introducing an entirely new analytic vocabulary at each stage.
The separation of Causal-Temporal subtypes may be particularly valuable in event-driven environments, where the learner does not see the dispatch mechanism that governs handler invocation. Treating Event as a distinct descriptive subtype enables classroom questions to explicitly surface hidden control, rather than treating all control as if it were user-authored Branch flow. The exploratory Year 5 probe in Section 4 suggests that early event-handling metaphors may shape pupils’ expectations about simultaneous handler activity, reinforcing the value of making hidden dispatch mechanisms explicit in classroom talk.
From a pedagogical perspective, PRIMM (sentance2019primm) provides a temporal sequence for classroom activity (Predict, Run, Investigate, Modify, Make). FLARE can complement this by providing a consistent descriptive vocabulary for the discussion that occurs within each phase. For example, during Predict and Run, teachers can stabilise elements and triggers; during Investigate, questions can focus on control movement (Causal-Temporal) and information movement (Communicative); and during Modify and Make, learners can use the same vocabulary to plan what a new element should receive, send, and share. This relationship is a hypothesis for future work rather than an established instructional effect.
@@
In curriculum design, the recursive model is compatible with a staged progression in which new abstraction scales appear once earlier bindings are secure. Early years may emphasise Sequential flow within small elements; later years can introduce Branch bindings and associated conditional reasoning, and then Event bindings and their hidden control mechanisms. Purpose can be introduced at any rung as an interpretive stance (”What is this for? Who relies on it?”) without adding another structural tier.
Beyond the system boundary: maintaining continuity with overlays
A recurring challenge in computing education is the transition from closed programs to open systems, where behaviour depends on environmental factors such as runtime scheduling, service contracts, network latency, permissions, partial failure, and platform policy. In these contexts, a purely code-centred structural description is often insufficient: not because the code loses structure, but because behaviour is increasingly co-produced by the interaction between code and its operating context.
FLARE v2 accommodates this transition by treating the element-and-binding model as a structural baseline and introducing overlays as continuity mechanisms above the system layer. Just as Event binding bridges the gap between authored control structures and hidden dispatch, overlays bridge the gap between a system’s internal composition and the external constraints that govern its execution. The learner does not discard structural analysis when moving to system-plus contexts; instead, they extend it.
Practically, this means the element-binding-element cycle remains the first pass for describing the code-side composition, while additional overlays make explicit the environmental forces acting upon that structure (for example, concurrency and scheduling assumptions, service-contract obligations, authority and trust boundaries, or failure semantics). This supports a spiral curriculum: earlier structural reasoning remains valid, but becomes insufficient on its own, and is extended rather than replaced.
@@ 8 Limitations and Future Work
The Year 5 probe reported in Section 4 is exploratory and requires systematic replication with controlled conditions. The analysis throughout is practitioner-led and interpretive, drawing on reflective re-reading of earlier classroom work rather than on a newly collected or systematically coded dataset.
Several specific claims require empirical investigation. The assertion that Event binding can often be approximated mechanically as Branch inside a loop but remains descriptively useful for classroom talk raises a testable question: do learners and teachers benefit from this distinction, or does it introduce unnecessary complexity? Similarly, the parallel between literacy progressions and the Compositional Ladder is offered as a potentially useful metaphor for teachers, but whether it reduces cognitive load in practice remains untested.
A further limitation is boundary sensitivity at system-plus. At services and ecosystems, the framework provides a structural baseline for code-side composition and typically needs complementary overlays for operational and socio-technical constraints. Future work should specify, test, and refine a small number of overlays that make these dependencies discussable in classrooms and analysable in research (for example, identity and aliasing for object-based reasoning; explicit concurrency and scheduling models; and service-contract and failure semantics for distributed interaction).
Future research should examine how the recursive binding cycle manifests in teacher and learner discourse. Studies might analyse classroom dialogue using the element-binding-element cycle, investigate how teachers use the Compositional Ladder to structure tasks and questions, or explore whether making hidden runtime control explicit supports more accurate notional machines for event-driven code. Implementation studies and evaluations of classroom uptake will be essential for assessing the practical value of FLARE v2.
@@ 9 Conclusion
This paper reframes FLARE from a four-tier descriptive hierarchy (Blocks, Segments, Relations, Macro) to a recursive element-binding-element cycle. The contribution is a compact descriptive scheme for discussing program meaning at increasing scale: bounded elements are described by what they Receive, Send, Effect, and Share; and elements bind through two dimensions, Causal-Temporal (Sequential, Branch, Event) and Communicative (Sends/Receives and Shares). The recursive cycle makes explicit how meaning can be composed across levels, rather than treating levels as a fixed classification.
The two-dimension scheme aligns with Pennington’s control-flow and data-flow distinction, while accommodating event-driven environments by treating hidden dispatch as a distinct descriptive subtype within Causal-Temporal binding. The scope of the claim remains as stated in Section 2: FLARE v2 is developed for imperative and event-driven environments typical of primary and lower-secondary curricula. At services and ecosystems, the framework provides a structural baseline for code-side composition and typically needs complementary overlays for operational and socio-technical constraints. This paper does not claim demonstrated coverage of object-oriented design reasoning, explicit concurrency models, distributed systems, or functional paradigms.
FLARE v2 is presented as a theoretical proposal grounded in classroom need. Its value should be judged by whether it proves generative for analysis, teacher questioning, and curriculum design, and by whether it supports targeted empirical investigation and refinement.
Across these boundaries, the intent is continuity: Event binding preserves structural reasoning where control transfer is mediated by hidden runtime dispatch, and overlays preserve that same reasoning where behaviour is further shaped by environmental constraints beyond the system boundary.
@@ 12 Conflict of Interest
This preprint is made available under the Creative Commons Attribution-NonCommercial-ShareAlike 4.0 International Licence (CC BY-NC-SA 4.0). This permits copying, distribution, and adaptation for non-commercial purposes, provided that appropriate credit is given and that derivative works are released under the same licence.
This licence applies only to the text, figures, and explanations contained in this paper. Practical implementations of the framework - including training materials, classroom scaffolds, exemplar resources, and software tools - are licensed separately and are not covered by this CC BY-NC-SA 4.0 licence.